OpenStudy (anonymous):

Let f(x)=(x^2)-sqrt(x). -show that x^2< or equal to f(x) for all x> or equal to 0 -use the inequality in the first part to to show that the integral of 1/f(x) from 1 to infinity converges. How do you start this?

OpenStudy (anonymous):

\(f(x)=x^2-\sqrt{x}\) domain-> [0,infty) for x>1,f(x)>0 for x=1, f(x)=0 for 0<x<1, f(x)<0

OpenStudy (anonymous):

\[\sqrt{x}=x^2-f(x)>0\] since square-root cannot be negative